OpenStudy (anonymous):

Match the values of x with the solution for y if y plus two and five−twelfths equals x. Match Term Definition x equals one−half A) y equals negative four and one−fourth x equals three and one−third B) y equals negative three and one−sixth x equals negative three−fourths C) y equals eleven−twelfths x equals negative one and five−sixths D) y equals negative one and eleven−twelfths

OpenStudy (anonymous):

y + 2 5/12 = x (change to improper fraction) y + 29/12 = x now if x = 1/2 y + 29/12 = 1/2 y = 1/2 - 29/12 y = 6/12 - 29/12 y = - 23/12 = -1 11/12 so if x = 1/2, then y = 1 1/12 (D) if x = 3 1/3(change to improper fraction) x = 10/3 y + 29/12 = 10/3 y = 10/3 - 29/12 y = 40/12 - 29/12 y = 11/12 so if x = 3 1/3, then y = 11/12 (C) if x = -3/4 y + 29/12 = -3/4 y = -3/4 - 29/12 y = - 9/12 - 29/12 y = - 38/12 = - 3 1/6 so if x = -3/4, then y = -3 1/6 (B) if x = - 1 5/6 (turn to improper fraction) x = - 11/6 y + 29/12 = - 11/6 y = -11/6 - 29/12 y = - 22/12 - 29/12 y = - 51/12 = - 4 1/4 so if x = -1 5/6, then y = -4 1/4 (A) any questions ?